OpenStudy (anonymous):

WILL MEDAL!!! I still need help please. okay guys i really need help explaining something. my little cousin is just learning algebra and i am trying to explain in the simplest terms how to find the value of x. I think all the ways ive show her are confusing her more so if you guys have any easier ways or tips to help i would be very appreciative and medal! thanks!!

OpenStudy (anonymous):

think of x as the number you are trying to find, you dont know that number, so x is used to replace that number you are trying to solve for

OpenStudy (anonymous):

Well one of her equations is What is the value of x when 2x + 3 = 3x – 4 ? so how can i explain to her how to do it?

OpenStudy (anonymous):

if you are trying to solve it, x=7

OpenStudy (anonymous):

tell her to replace the x with 7

OpenStudy (anonymous):

so it would be 2 x 7 + 3 = 3 x 7 - 4

OpenStudy (anonymous):

alright but i want to explain how to find 7 like how did you come up with x=7 im trying to teach her but im terrible at explaining

OpenStudy (anonymous):

simplifys into 14 + 3 = 21 - 4, so this would equal 17 = 17

OpenStudy (anonymous):

wait, one sec

OpenStudy (anonymous):

do u know how to solve this?

OpenStudy (anonymous):

or do u want me to tell u why u do each step?

OpenStudy (anonymous):

no i know how. Im trying to teach her how to find the value of x. What is the value of x when 2x + 3 = 3x – 4 ? so i want to explain to her how to find what what x equals. like when you said x = 7 how did you find out that x=7? what steps did you take to find that? im trying to explain that to her

OpenStudy (anonymous):

im not trying to solve it. Im trying to show her how to find the value of x thats it. This is an example question.

OpenStudy (anonymous):

i dont really know why lol, i do see your side haha, its harder trying to explain then solving

OpenStudy (anonymous):

Exactly!! thats my problem. I can get the answer that x=7 but i cant explain how to get it. so i need help explaining it to her in general so she can apply it to all her algebra problems like this

OpenStudy (anonymous):

@heretohelpalways any ideas??

Vocaloid (vocaloid):

2x + 3 = 3x – 4 1. subtract 2x from each side 3 = x - 4 2. add 4 to both sides 7 = x boom

OpenStudy (anonymous):

Thanks for that. Do you have an easy way to explain how to do this on all types of problems? like how to find x on other problems as well?

Vocaloid (vocaloid):

basically, we just want to manipulate the equation to get x by itself you can add/subtract the same thing from both sides of the equation, until we have x by itself on one side

Vocaloid (vocaloid):

the best way to explain it, is that "x" represents an unknown number, and we want to get x by itself to see what it's equal to we can add/subtract/multiply/divide.etc. as long as we do the same thing to both sides of the equal sign

OpenStudy (ddos_dragon):

Inverse operations using the same number always equal 0. 3* 5 / 3 You can take those threes out. To find X, you can use the inverse operation "trick" To eliminate operations and the number that comes with it.

OpenStudy (ddos_dragon):

Medal Vocaloid, if you already haven't and close the question. Have a nice day!